Zoogle has the following selected data ($ in millions): (Round your answers to 2 decimal place. Enter your answers in million2. Calculate the cash return on assets. ($ in millions) Cash Return on Assets Zoogle 3. Calculate the cash flow to sales rati($ in millions) Asset Turnover times Zoogle
Answer #1

1. Return on assets = Net income / Average assets

= 6440/(23768+32497)/2 = 22.89%

2. Cash return on assets = operating cash flows / average assets

= 9308/(23768+32497)/2 = 33.09%

3. Cash flow to sales = operating cash flow / sales

= 9308/22851 = 0.41

4. Assets turnover = sales / average assets

= 22851/(23768+32497)/2 = 0.81 times